A standard solution of iron was made by weighing 0.075 g of Fe(NH4)2(SO4)2 6(H2O) in 250 mL. Aliquots of this standard solution (see below) were transferred to a 100 mL volumetric flask, pH adjusted with citrate, and reacted with hydroquinone and o-phenanthroline and diluted to volume (100 mL). The absorbance was measured in a 1.0-cm cell using a Genesys 20 Spectrophotometer at 508 nm:

Aliquot of standard solution Absorbance
1.0 mL 0.079
2.0 mL 0.163
5.0 mL 0.413
10.0 mL 0.842

A pill weighing 0.410 g was dissolved in 25 mL HCl and diluted to 100 mL. Of this solution, a 10.0 mL aliquot was transferred to a clean 100-mL volumetric flask and diluted to volume. Finally, a 10.0-mL aliquot of this dilute solution was added to a clean 100-mL volumetric flask with 6 mL citrate, 2.00 mL of hydroquinone, and 3.00 mL o-phenanthroline and diluted to the mark (100.0 mL). After standing for 10 minutes, the absorbance was read 3 times in a 1.0 cm cell, and the average absorbance was 0.226.

Use these data to report:

(1) the molar absorptivity of the iron o-phenanthroline complex at 508nm, and

(2) the mass of iron in the pill sample in mg.
Answer #1

Standard solution

molar concentration of stock solution = 0.075 g/392.13 g/mol x 0.250 L = 7.65 x 10^-4 M

Aliquots diluted to 100 ml

Using M2 = M1V1/V2

with M2 is final concentration, M1 is initial stock solution concentration, V1 and V2 are initial and final solution volumes

Volume of stock (ml)      Final concentration (M)

                 1                           7.65 x 10^-6

                 2                           1.53 x 10^-5

                 5                           3.82 x 10^-5

                10                          7.65 x 10^-5

Plot Absorbance on y-axis and concentration on x-axis.

(1) slope = molar absorptivity of iron-phenanthroline complex = (0.842 - 0.079)/(7.65 x 10^-5 - 7.65 x 10^-6) = 11082 M-1.cm-1

(2) Unknown

Absorbance = 0.226

concentration of complex in final diluted solution = 0.226 x 7.65 x 10^-5/0.842 = 2.053 x 10^-5 M

concentration of complex in original 100 ml solution of pill = 2.053 x 10^-5 x 100 = 2.053 x 10^-3 M

mass of iron in the pill = 2.053 x 10^-3 M x 100 ml x 55.84 g/mol = 11.4 mg